# Implicit Bias in Matrix Factorization and its Explicit Realization in a New Architecture

### v1

@@ 3.2 Numerical Experiments on Matrix Factorization
Next, we investigate the impact of initialization. We fix the step-size at $\eta=10^{-2}$ and evaluate the effect of varying $\xi$. We observe a correlation between the implicit bias of the BM factorization and $\xi$, which determines the initial distance from the origin. Initializing closer to the origin in the classical BM factorization yields solutions with a faster spectral decay. Notably, the UDU factorization demonstrates a strong implicit bias toward truly low-rank solutions, regardless of the choice of $\eta$ or $\xi$.
We provide additional experiments on matrix factorization problems in the Appendices. Specifically, Section A.1 considers the matrix completion problem with noisy measurements, where $b$ is perturbed with Gaussian noise. The results remain consistent with the noiseless case: the UDU model exhibits an implicit bias toward truly low-rank solutions, while the classical BM factorization yields approximately low-rank solutions, with spectral decay strongly influenced by initialization.
In Section A.2, we present the evolution of the factors $U$ and $D$ over the iterations. Interestingly, the UDU model produces rank-revealing solutions, with the columns of $U$ converging to zero in certain directions. Specifically, $U$ tends to grow along certain directions, while the rescaling induced by projection onto the bounded constraint shrinks other coordinates. This behavior aligns with the mechanism of the power method and offers insights into its connection with divergent forces.
Additionally, we present numerical experiments on a matrix sensing problem arising in phase retrieval image recovery in Section A.3. As before, the UDU framework consistently promotes low-rank solutions, and this structural bias significantly enhances the quality of the recovered image, as demonstrated by our results.
4 Feedforward Neural Networks with Diagonal Hidden Layers
This section extends our approach to neural networks. Consider a dataset comprising $n$ data points $(\mathbf{x}_{i},\mathbf{y}_{i})\in\mathbb{R}^{d}\times\mathbb{R}^{c}$. We first define a three-layer neural network defined as
$\phi(\mathbf{x}):=\sum_{j=1}^{m}\mathbf{v}_{j}w_{j}\mathbf{u}_{j}^{\top}\mathbf{x}~{}~{}\approx~{}~{}\mathbf{y}.$ (6)
The first and third layers are fully connected, and the middle is a diagonal layer, as illustrated in Figure 2. Drawing parallels between our matrix factorization model in (4) and neural network training, we impose Euclidean norm constraints on the weights of the fully connected layers. Under these conditions, the training problem can be formulated as follows:
$\min_{\begin{subarray}{c}\mathbf{u}_{j}\in\mathbb{R}^{d}\\ \mathbf{v}_{j}\in\mathbb{R}^{c}\\ w_{j}\in\mathbb{R}\end{subarray}}$ $\frac{1}{2n}\sum_{i=1}^{n}\|\sum_{j=1}^{m}\mathbf{v}_{j}\,w_{j}\,\mathbf{u}_{j}^{\top}\mathbf{x}_{i}-\mathbf{y}_{i}\|_{2}^{2}$ (7)
$\mathrm{subj.to}$ $\sum_{j=1}^{m}\|\mathbf{u}_{j}\|_{2}^{2}\leq 1,~{}~{}\sum_{j=1}^{m}\|\mathbf{v}_{j}\|_{2}^{2}\leq 1,~{}~{}\text{and }~{}~{}w_{j}\geq 0;\quad\text{for all}~{}~{}j=1,\ldots,m.$
The norm constraints in our training problem can be interpreted as a stronger form of weight decay, one of the most commonly used regularization techniques in neural networks, which lends further justification to our formulation. We refer to this neural network structure as UDV.
@@ 4 Feedforward Neural Networks with Diagonal Hidden Layers
4.1 Numerical Experiments on Neural Networks
In this section, we test the proposed UDV framework on regression and classification tasks, comparing it with fully connected two-layer neural networks (denoted as UV in the subsequent text) using both linear and ReLU activation functions. This comparison is fair in terms of computational cost, as the cost incurred by the diagonal layer —which can also be viewed as a parameterized linear activation function— is negligible. We demonstrate a strong empirical bias toward low-rank solutions in all our experiments. We also present a proof-of-concept use case of this strong bias, combined with an SVD-based pruning strategy, to produce compact networks.
@@ 4.1.1 Implementation Details
Computing environment. All classification tasks were conducted on an NVIDIA A100 GPU with four cores of the AMD Epyc 7742 processor, while regression tasks were conducted on a single core of an Intel Xeon Gold 6132 processor. We used Python 3.9.5 and PyTorch 2.0.1.
Datasets. We used two datasets for the regression tasks: House Prices - Advanced Regression Techniques (HPART) (Anna Montoya, 2016) and New York City Taxi Trip Duration (NYCTTD) (Risdal, 2017). We allocated 80% of the data for the training and reserved the remaining 20% for validation. Following (Huang, 2003), we set the number of hidden neurons in the diagonal layer $\smash{m=\texttt{round}\big{(}\sqrt{(c+2)d}+2\sqrt{d/(c+2)}\big{)}}$. This results in a network structure ($d$-$m$-$c$) of 79-26-1 for HPART, and 12-10-1 for NYCTTD.
For classification tasks, we used the normalized MNIST dataset (LeCun et al., 2010). We applied transfer learning by replacing the classifier layers of three advanced neural networks –MaxViT-T (Tu et al., 2022), EfficientNet-B0 (Tan & Le, 2019), and RegNetX-32GF (Radosavovic et al., 2020)– with UDV, while using pre-trained weights from ImageNet-1K (Deng et al., 2009). Specifically, we retained all layers up to the first fully connected layer of the classifier and replaced the subsequent layers. The number of hidden neurons in the diagonal layer was set to as $m=\texttt{floor}(\tfrac{2}{3}d)$. This results in a UDV network structure ($d$-$m$-$c$) of 512-341-10 for MaxViT-T, 1280-853-10 for EfficientNet-B0, and 2520-1680-10 for RegNetX-32GF.
@@
Optimization methods. We tested the results using four different optimization algorithms for training: Adam (Kingma & Ba, 2014), Mini-Batch Gradient Descent (MBGD) (LeCun et al., 2002), NAdam (Dozat, 2016), and Mini-Batch Gradient Descent with Momentum (MBGDM) (Sutskever et al., 2013). For classification, we used different batch sizes for different models: 128 for MaxViT-T and RegNetX-32GF, and 384 for EfficientNet-B0.
We tuned the step size for all models and optimization algorithms: For regression tasks, we tested step sizes $10^{-4}$, $10^{-3}$, $10^{-2}$, $10^{-1}$, $1$, $2$, $3$. Larger step sizes ($1$, $2$, $3$) were often excluded for the UV model due to divergence. In classification experiments, we tested learning rates $10^{-6}$, $10^{-5}$, $10^{-4}$, $10^{-3}$, $10^{-2}$, $10^{-1}$, $1$ with Adam and NAdam; and we tested $10^{-3}$, $10^{-2}$, $10^{-1}$, $1$, $2$, $3$, $5$ with MBGD and MBGDM.
Training Procedure. UV and UDV models were initialized identically, with the diagonal elements of $D$ initialized using Kaiming Uniform Initialization, consistent with the default initialization for fully connected layers in PyTorch. The models were trained for 200 epochs on HPART, 50 epochs on NYCTTD, and 70 epochs on MNIST; and results were averaged over 1000 random seeds for HPART, 100 for NYCTTD, and 1 for MNIST to ensure robustness. The validation loss, used as a generalization metric in regression, was averaged over the final 20 epochs for the HPART dataset and the final 5 epochs for the NYCTTD dataset. Similarly, validation accuracy for classification tasks was averaged over the last 5 epochs to ensure stability in the reported values.
Repository. The code is available at https://github.com/Titanium-H/UDV
4.1.2 Low-rank Bias in Neural Network Training
Table 1 presents the validation loss (for regression) or validation accuracy (for classification) of the UDV model compared to the classical UV model with linear and ReLU activation functions. For each configuration (dataset and model architecture), the results are obtained by selecting the best algorithm and learning rate pair. Moreover, Figure 3 illustrates the singular value spectrum of the solutions corresponding to each entry in these tables. We focus on the singular values from the $U$ and $UD$ layers, as they generate the primary data representation, while omitting the $V$ layer, which serves as the feature selection layer and is a tall matrix by definition, given that $c\ll m$ in most cases. Collectively, these results show that the UDV framework achieves competitive prediction accuracy while exhibiting a strong implicit bias toward low-rank solutions, as indicated by the faster decay in the singular value spectrum.
@@ 4.1.2 Low-rank Bias in Neural Network Training
Table 1: Model performance using different models. M, E, and R represent the transferred models MaxVit-T, EfficientNet-B0, and RegNetX-32GF, respectively.
Tasks Regression (Test Loss) Classification (Test Accuracy)
Dataset HPART NYCTTD MNIST
UDV
$1.304\times 10^{-3}$
Adam: $10^{-3}$
$5.248\times 10^{-6}$
NAdam: $10^{-4}$
M: 99.67%
MBGDM: $10^{-2}$
E: 99.63%
MBGDM: $10^{-1}$
R: 99.74%
MBGDM: $10^{-2}$
UV
$1.333\times 10^{-3}$
Adam: $10^{-3}$
$5.251\times 10^{-6}$
Adam: $10^{-3}$
M: 99.69%
MBGDM: $10^{-2}$
E: 99.60%
Adam: $10^{-3}$
R: 99.66%
MBGD: $10^{0}$
UV-ReLU
$1.167\times 10^{-3}$
Adam: $10^{-3}$
$5.323\times 10^{-6}$
NAdam: $10^{-3}$
M: 99.68%
NAdam: $10^{-4}$
E: 99.68%
MBGDM: $10^{-1}$
R: 99.73%
MBGD: $10^{0}$
Figure 3: Singular value spectrum corresponding to the solutions reported in Table 1.
4.1.3 Reducing Network Size with SVD-based Pruning

### v2

@@ 2.2.1 Matrix Completion
Next, we examine the impact of initialization. We fix the step-size at $\eta=10^{-2}$ and evaluate the effect of $\xi$. In the classical BM factorization, initializing closer to the origin leads to solutions with faster spectral decay, consistent with the observations of Gunasekar et al. (2017). In contrast, the proposed UDU factorization exhibits a pronounced implicit bias toward exactly low-rank solutions, independent of the choice of $\xi$.
We also conducted the same experiments under Gaussian noise in Section A.1, and the results are qualitatively consistent. The UDU model continues to exhibit an implicit bias toward truly low-rank solutions, while the classical method gets approximately low-rank solutions. Exact rank-3 recovery is no longer possible in the noisy setting, but the UDU factorization still generates a truly low-rank solution, at a slightly higher rank.
2.2.2 Fourier Ptychography
We consider a phase retrieval problem that arises in Fourier ptychography (FP), an imaging technique that reconstructs the phase of a complex-valued image (the sample’s transmission function) from multiple Fourier measurements under illuminations from different angles. We use real measurements acquired from a functional FP system applied to a human blood cell sample, obtained from Horstmeyer et al. (2015). The target is a $160\times 160$ complex-valued image, vectorized into $d=160^{2}=25^{\prime}600$ dimensions and represented by a vector $x_{\natural}\in\mathbb{C}^{d}$.
The measurements are modeled as $b_{i}\approx|\langle a_{i},x_{\natural}\rangle|^{2}$ for $i=1,\dots,n$, where the approximation accounts for experimental noise from sensor limitations, illumination variability, and other imperfections. The sensing vectors $a_{i}\in\mathbb{C}^{d}$ are derived from windowed Fourier transforms under different illumination patterns. The dataset consists of $29$ illumination patterns, each recorded as an $80\times 80$ intensity image, yielding a total of $n=185^{\prime}600$ measurements.
FP is a nonconvex quadratic inverse problem, which admits a lifted formulation as a PSD matrix recovery problem with a rank-1 constraint. The lifted formulation corresponds to recovering a $25^{\prime}600\times 25^{\prime}600$ complex PSD matrix from the measurements, which makes solving at full-rank factorization (with $r=d$) computationally infeasible. Instead, we set the factorization rank to $r=20$ and compare the proposed UDU formulation with the standard BM factorization. Both methods are run for $10^{4}$ iterations with step size $0.1$, initialized from the same random $U_{0}$ with i.i.d. Gaussian complex entries normalized in Frobenius norm. For UDU, we set $D_{0}=I$ and $\alpha=1$. Once solved, we recover the image estimate by extracting the dominant singular vector from the ($d\times r$) dimensional factors.
The results are shown in Figure 2. The UDU factorization successfully constructs the image, whereas BM produces a heavily corrupted reconstruction. An inspection of the singular value spectrum reveals that UDU effectively converges to a rank-1 solution, while BM yields a spectrum with slow decay, explaining the difference in reconstruction quality. Importantly, UDU achieves this from a random initialization, in contrast to existing nonconvex phase retrieval methods that typically require carefully designed spectral initialization procedures (Candes et al., 2015; Bian et al., 2015).
Figure 2: [Left] Singular value spectra of the solutions obtained by BM and UDU after $10^{4}$ iterations with step size $0.1$. [Middle] The corresponding reconstructed image from the BM factorization exhibits artifacts. [Right] The UDU factorization produces a clean recovery.
Additional details are provided in Section A.2. There, we show reconstructions at intermediate iterations, and experiments on smaller synthetic phase retrieval problems using full-rank factorizations ($r=d$). In all cases, the UDU framework promotes low-rank solutions, and this structural bias improves phase recovery.
2.3 Theoretical Insights into the Inner Workings and Implicit Bias
A fixed-point analysis of the proposed method provides valuable insights into its inner workings.
Lemma 1 (Fixed-point characterization).
Define the update variables before projection as $\bar{U}=U-2\eta\nabla f(X)UD$ and $\bar{D}=D-\eta U^{\top}\nabla f(X)U$, with $X=UDU^{\top}\!$. Suppose $(U,D)$ is a fixed point of the algorithm in (3), and let $u_{j}$ denote the $j^{\text{th}}$ column of $U$ and $\lambda_{j}$ the $j^{\text{th}}$ diagonal entry of $D$. Then,
  1. (a)
    If $\|\bar{U}\|_{F}\leq\alpha$, then $\nabla f(X)u_{j}\lambda_{j}=0$ for all $j$.
  2. (b)
    If $\|\bar{U}\|_{F}>\alpha$, then there exists some $\beta>0$ such that $\nabla f(X)u_{j}\lambda_{j}=-\beta u_{j}$ for all $j$.
  3. (c)
    If $\lambda_{j}=0$, then $\smash{u_{j}^{\top}\nabla f(X)u_{j}\geq 0}$.
  4. (d)
    If $\lambda_{j}>0$, then $\smash{u_{j}^{\top}\nabla f(X)u_{j}=0}$.
The proof is given in Section A.4.
At first glance, condition $(b)$ might suggest that choosing a smaller $\alpha$ leads to fixed points where the columns of $U$ align with the negative eigenvectors of $\nabla f(X)$. However, the following proposition shows that no valid fixed point can satisfy $\|\bar{U}\|_{F}>\alpha$.
Proposition 1 (Exclusion of spurious fixed points).
There are no fixed points $(U,D)$ of the algorithm in (3) such that $\|\bar{U}\|_{F}>\alpha$. As a consequence, all valid fixed points satisfy $\|\bar{U}\|_{F}\leq\alpha$, and the fixed point structure coincides with that of BM factorization (after the change of variables $U\mapsto UD^{1/2}$). Thus, adding the diagonal factor $D$ and bounding $U$ does not create new fixed points for $X$.
Proof.
Suppose $\|\bar{U}\|_{F}>\alpha$ and $\lambda_{j}>0$. Then, by combining Lemma 1(b)\,(b) and $(d)$, we get
$0\overset{(d)}{=}u_{j}^{\top}\nabla f(X)u_{j}\overset{(b)}{=}-\tfrac{\beta}{\lambda_{j}}\|u_{j}\|^{2}.$
which forces $u_{j}=0$. If instead $\lambda_{j}=0$, then $(b)$ again implies $u_{j}=0$. Thus in both cases the columns of $U$ must vanish, leading to $U=0$, and hence $\bar{U}=0$, which contradicts the assumption $\|\bar{U}\|_{F}>\alpha$. Hence, for any fixed point of the algorithm, the constraint is inactive and $\|\bar{U}\|_{F}\leq\alpha$.
It remains to characterize these fixed points. From Lemma 1(a)(a), any fixed point satisfies $\nabla f(X)u_{j}\lambda_{j}=0$ for all $j$. Define $\hat{u}_{j}:=u_{j}\sqrt{\lambda_{j}}$ and let $\hat{U}=[\hat{u}_{1},\ldots,\hat{u}_{r}]$. Then $\hat{U}\hat{U}^{\top}=UDU^{\top}=X$ and $\nabla f(X)\hat{u}_{j}=0$ for all $j$, which coincides with the fixed-point condition of the BM factorization. Hence, the fixed points of our method correspond exactly to those of BM.
Conversely, let $\hat{U}$ be a stationary point of the BM factorization, so $X=\hat{U}\hat{U}^{\top}$ and $\nabla f(X)\hat{U}=0$. For any $\alpha>0$, set $U=(\alpha/\|\hat{U}\|_{F})\hat{U}$ and $D=(\|\hat{U}\|_{F}^{2}/\alpha^{2})I$. Then $\|U\|_{F}=\alpha$, $UDU^{\top}=X$, and $\nabla f(X)UD=0$. Thus, every stationary point of BM corresponds to a stationary point of UDU, and vice versa, at the level of $X=UDU^{\top}=\hat{U}\hat{U}^{\top}$. ∎
Although the fixed points of UDU coincide with those of BM, the above analysis still provides insight into the low-rank bias of the algorithm. In particular, when $U$ grows such that $\lVert\bar{U}\rVert\geq\alpha$, the updates tend to amplify directions aligned with the negative eigenvectors of $\nabla f(X)$. Concretely, we can express the update rules in terms of the columns of $u_{j}$ and $\lambda_{j}$ as $\bar{u}_{j}=u_{j}-2\eta\nabla f(X)u_{j}\lambda_{i}$ and $\smash{\bar{\lambda}_{j}=\lambda_{j}-\eta u_{j}^{\top}\nabla f(X)u_{j}}$. These expressions show that both $\bar{u}_{j}$ and $\bar{\lambda}_{j}$ tend to grow in the directions aligned with the negative eigenvectors of $\nabla f(X)$, even though, by Lemma 1(d)(d), no fixed point with $\lambda_{j}>0$ is possible unless $u_{j}^{\top}\nabla f(X)u_{j}=0$. This tension suggests that the dynamics either suppress some columns of $U$ (via projection) or drive $X$ toward configurations where $\nabla f(X)u_{j}=0$, thereby reinforcing low-rank structure along the path to convergence.
It is worth noting, that this discussion is a simplified perspective aimed at gaining intuition. In reality, the alignment or shrinking of the columns of $U$ and the minimization of $f(X)$ along these directions occur simultaneously and interact in a complex manner. Nevertheless, we can clearly observe these effects in our numerical experiments. In Section A.3, we present the evolution of the column norms of $U$ and the diagonal entries of $D$ over the iterations in our matrix completion experiment. Our results show that initially, a few specific columns of $U$ grow, pushing the others numerically to zero. Once $f(X)$ is effectively minimized with respect to these initial columns, additional columns are identified and begin to grow. Eventually, the algorithm converges to a low-rank solution, with $UDU^{\top}$ becoming rank-revealing since only a few columns of $U$ remain nonzero. We further observe that these nonzero columns are orthogonal, highlighting how the algorithm’s tendency to align $u_{j}$ with negative eigenvectors of $\nabla f(X)$ implicitly induces structured low-rank solutions along the path.
3 Neural Networks with Diagonal Hidden Layers
@@
This section extends our approach to neural networks. We introduce a three-layer neural network, with the first and third layers fully connected and the middle layer diagonal, as illustrated in Figure 3.
Let $\{(\mathbf{x}_{i},\mathbf{y}_{i})\}_{i=1}^{n}\subset\mathbb{R}^{d}\times\mathbb{R}^{c}$ denote a dataset of size $n$. The resulting network is given by
$\phi(\mathbf{x}):=\sum\nolimits_{j=1}^{m}\mathbf{v}_{j}w_{j}\mathbf{u}_{j}^{\top}\mathbf{x}$ (4)
with the goal of approximating $\phi(\mathbf{x}_{i})\approx\mathbf{y}_{i}$. Drawing parallels between our matrix factorization model in (2) and neural network training, we impose Euclidean norm constraints on the weights of the fully connected layers. Under these conditions, the training problem can be formulated as follows:
$\min_{\mathbf{u}_{j},w_{j},\mathbf{v}_{j}}$ $\frac{1}{2n}\sum\nolimits_{i=1}^{n}\|\sum\nolimits_{j=1}^{m}\mathbf{v}_{j}\,w_{j}\,\mathbf{u}_{j}^{\top}\mathbf{x}_{i}-\mathbf{y}_{i}\|_{2}^{2}$ (5)
$~\mathrm{s.t.}$ $\sum\nolimits_{j=1}^{m}\|\mathbf{u}_{j}\|_{2}^{2}\leq 1,~~\sum\nolimits_{j=1}^{m}\|\mathbf{v}_{j}\|_{2}^{2}\leq 1,$
$\text{and }~~w_{j}\geq 0;\quad\text{for all}~~j=1,\ldots,m.$
The norm constraints can be interpreted as a stronger form of weight decay, a standard regularization in neural networks, which lends further justification to our formulation. We refer to this neural network structure as UDV.
3.1 Experiments on Neural Networks
In this section, we test the UDV structure on regression and classification tasks, comparing it with fully connected two-layer neural networks (denoted as UV in the subsequent text). This comparison is fair in terms of computational cost since the overhead introduced by the diagonal layer, which can also be regarded as a parameterized linear activation function, is negligible. To further ensure fairness in comparison, no additional fine-tuning was introduced for the proposed method.
We observe a strong empirical bias toward low-rank solutions in all our experiments with the UDV structure. We also present a proof-of-concept use case of this strong bias, combined with an SVD-based pruning strategy, to produce compact networks.
@@ 3.1.1 Implementation Details
Datasets. We used two datasets for the regression tasks: House Prices - Advanced Regression Techniques (HPART) (Anna Montoya, 2016) and New York City Taxi Trip Duration (NYCTTD) (Risdal, 2017). We allocated 80% of the data for the training and reserved the remaining 20% for test. Following (Huang, 2003), we set the number of hidden neurons in the diagonal layer $\smash{m=\texttt{round}\big(\sqrt{(c+2)d}+2\sqrt{d/(c+2)}\big)}$. This results in a network structure ($d$-$m$-$c$) of 79-26-1 for HPART, and 12-10-1 for NYCTTD.
For classification tasks, we used the CIFAR-10 and CIFAR-100 dataset (Krizhevsky, 2009). We applied transfer learning to three modern high-performing architectures: MaxViT-T (Tu et al., 2022), EfficientNet-B0 (Tan & Le, 2019), and RegNetX-32GF (Radosavovic et al., 2020). The original classification head (the final linear layer) of each model was replaced with two linear layers, forming the backbone classifier (UV). By inserting a diagonal layer between them, the classifier was transformed into the proposed UDV structure. The new linear layers were initialized with a truncated normal distribution, and the diagonal layer was initialized as an identity matrix. The vector dimension followed that of the preceding layer, resulting in a UDV classifier structure ($d$-$m$-$c$), where $c$ denotes the number of classes (10 for CIFAR-10 and 100 for CIFAR-100). Specifically, the structures are 512-512-$c$ for MaxViT-T, 1280-1280-$c$ for EfficientNet-B0, and 2520-2520-$c$ for RegNetX-32GF.
@@
Optimization methods. We tested the results using four different optimization algorithms for training: Adam (Kingma & Ba, 2014), Mini-Batch Gradient Descent (MBGD) (LeCun et al., 2002), NAdam (Dozat, 2016), and Mini-Batch Gradient Descent with Momentum (MBGDM) (Sutskever et al., 2013). We tuned the learning rates (LRs, the term learning rates is used instead of step sizes for clarity in the machine learning context) for all models and optimization algorithms. In addition, a cosine annealing learning rate scheduler was applied to stabilize training and improve convergence.
Training Procedure and Metrics. The models were trained for 200 epochs on HPART, 50 epochs on NYCTTD, and 100 epochs on CIFAR-10 and CIFAR-100. Test loss for regression was averaged over the last 20 and 5 epochs, respectively, whereas test accuracy for classification was averaged over the final 3 epochs.
3.1.2 Low-rank Bias in Neural Networks
Table 1 presents the test loss (for regression) or test accuracy (for classification) of the UDV model compared to the classical UV model. For each configuration (dataset and model architecture), the results are obtained by selecting the best algorithm and learning rate pair. On CIFAR-100, the UDV model underperforms UV by about 1% in test accuracy. We note, however, that UDV achieves its best performance earlier in training (around 90 epochs), suggesting that early stopping may close this gap.
@@ 3.1.2 Low-rank Bias in Neural Networks
Table 1: Comparison of models with their best performance. M, E, and R represent the transferred models MaxVit-T, EfficientNet-B0, and RegNetX-32GF, respectively.
Tasks Regression (Test Loss) Classification (Test Accuracy %)
Dataset HPART NYCTTD CIFAR-10 CIFAR-100
UDV $1.304\times 10^{-3}$ $5.248\times 10^{-6}$
M: 99.17
E: 97.87
R: 98.67
M: 86.18
E: 91.63
R: 89.79
UV $1.333\times 10^{-3}$ $5.251\times 10^{-6}$
M: 99.11
E: 97.99
R: 98.65
M: 87.75
E: 92.03
R: 90.33
Figure 5 (top row) shows the singular value spectra for RegNetX-32GF on CIFAR-100 under different optimization algorithms. We report the spectra for the $U$ (for the classical model) and $UD$ (for the proposed model) layers, which generate the main data representation, and omit the V layers since it primarily acts as a feature selector and is typically tall ($c\ll m$). Collectively, the results show that the UDV framework matches the predictive accuracy of UV while exhibiting a clear implicit bias toward low-rank solutions, as reflected in the faster spectral decay.
Figure 4 presents similar results for the HPART regression task. The findings are consistent, as UDV induces a strong low-rank bias while maintaining predictive performance comparable to UV.
Figure 4: Singular value spectra of the solutions obtained with UV and UDV formulations on the HPART regression task under different optimization algorithms.
3.1.3 Reducing Network Size via SVD-based Pruning